OpenStudy (anonymous):

2x-3xy=4-3y solve for y

OpenStudy (anonymous):

3xy-3y=2x-4 y(3x-3)=2x-4 then y=(2x-4)/(3x-3) so x can't be 1 if x=1 we have 2-3y=4-3y which is obviously false so you get that for each x except x=1, y=(2x-4)/(3x-3)

OpenStudy (anonymous):

how did u get from the original equation 2x-3xy=4-3y to 3xy-3y=2x-4

OpenStudy (anonymous):

2x-3xy=4-3y multiply by -1 -2x+3xy=-4+3y then substract -2x from both sides 3xy=-4+3y+2x then substract 3y from both sides 3xy-3y=-4+2x 3xy-3y=2x-4 addition and multiplcation are associative and commutative: that is a+b=b+a (a+b)+c=a+(b+c) also a*b=b*a and (a*b)*c=a*(b*c)